# MGMCpaper.bib

### v2

@@ 3.2 Random smoothers
A central ingredient of Alg. 1 is the $\mathrm{RandomSmoother}()$ procedure in lines 7 and 16, which can be seen as a generalisation of the Gibbs smoother. All random smoothers considered here are based on splitting a general symmetric matrix $A$ into two parts as $A=:M-N$, where the matrix $M$ defines the splitting. The generic update procedure is written down in Alg. 2 where here and in the following we assume that the input $\theta$ is drawn from a multivariate normal distribution. As the following lemma shows, it is always possible to symmetrise a given random smoother defined by the splitting $A=M-N$.
Algorithm 2 Random Smoother for updating $\theta\mapsto\theta^{\prime}$ based on the matrix splitting $A=M-N$
1:procedure $\mathrm{RandomSmoother}$($A$, $M$, $f$, $\theta$)
@@ 3.3 Relationship to deterministic smoothers
would result in a deterministic smoother defined by the splitting $A=M-N$ (see e.g. [saad2003iterative, section 4.2.2]). Under certain conditions on the splitting matrix $M$, repeated applications of the update in (3.4) can be used to iteratively solve the linear system $Au=f$. For example, when splitting the symmetric matrix $A$ into its diagonal $D$, strict lower triangular part $L$ and strict upper diagonal part $L^{\top}$
@@
$M$ $=M^{\mathrm{SOR}}=\frac{1}{\omega}D+L,$ $N$ $=N^{\mathrm{SOR}}=\frac{\omega-1}{\omega}D-L^{\top}\qquad\text{with $\omega>0$}.$ (3.6)
As discussed extensively in [Fox.C_Parker_2017_AcceleratedGibbsSampling], there is a close relationship between classical smoothers for the solution of the linear system $Au=f$ and what we call random smoothers for sampling from the multivariante normal distribution $\mathcal{N}(\mu,A^{-1})$ with $\mu=A^{-1}f$ in Section 3.2. For example, the random pendant of the deterministic SOR smoother in (3.6) with $\omega=1$ is the Gibbs-sampler (see [Fox.C_Parker_2017_AcceleratedGibbsSampling, Table 2]), which we will refer to as the Gibbs-smoother from now on to make the connection more explicit.
@@ 3.4 Random smoother with low-rank update for Bayesian inference
$\theta^{\prime}_{\ell}$ $=\theta_{\ell}^{*}-B_{\ell}^{*}(B_{\ell}^{\top}\theta_{\ell}^{*})$
@@
$B_{\ell}^{*}=(\omega^{-1}D_{\ell}+L_{\ell})^{-1}B_{\ell}\left(\Gamma+B_{\ell}^{\top}\left(\omega^{-1}D_{\ell}+L_{\ell}\right)^{-1}B_{\ell}\right)^{-1},$ (3.13)
which can be precomputed once at the beginning of the simulation. An equivalent backward Gibbs sampler with low rank correction can be obtained by swapping $L_{\ell}\leftrightarrow L_{\ell}^{\top}$ in (3.12) and (3.13). Putting everything together we arrive at Alg. 3. Obviously, $C_{\ell}$ and $B_{\ell}^{*}$ can be precomputed once and used in all subsequent calls, so that the setup costs are amortised for large numbers of samples.
@@
8:     Compute $B_{\ell}^{*}=C_{\ell}\left(\Gamma+B_{\ell}^{\top}C_{\ell}\right)^{-1}$
@@
$\theta^{\prime}_{\ell}=\theta_{\ell}^{*}-B_{\ell}^{*}(B_{\ell}^{\top}\theta_{\ell}^{*})$ (3.14)
@@ 3.5 Coarse level sampler
On the coarsest level we have to make a choice for the $\mathrm{CoarseSampler()}$ procedure that generates a new sample $\theta^{\prime}_{0}$ in line 3 of Alg. 1. Using Alg. 5, which consists of repeated applications of the random smoother in Alg. 2, corresponds to the approach in [goodman1989multigrid]. It turns out that this choice is sufficient to guarantee that Alg. 1 leaves the target distribution invariant (see Section 4.2).
@@ 3.6 Cost analysis
In this section, we derive an upper bound on the cost of one MGMC update in Alg. 1 as a function of the problem size $n_{L}$ and the rank $\beta$ of the measurement operator $B_{L}$. We analyse this cost under the assumption that the matrices $A_{\ell}$, $I_{\ell-1}^{\ell}$ and $I_{\ell}^{\ell-1}$ are sparse, with some upper bound on the number of entries per row, independent of the problem size. For the local basis functions in finite element spaces considered here this is naturally the case. Consequently, any matrix-vector products with $A_{\ell}$ or $I_{\ell-1}^{\ell}$ incur a cost of $\mathcal{O}(n_{\ell})$, in the same way as any tridiagonal solve $(\omega^{-1}D_{\ell}+L_{\ell})\theta_{\ell}=r_{\ell}$ or $(\omega^{-1}D_{\ell}+L_{\ell}^{\top})\theta_{\ell}=r_{\ell}$ for a single right-hand side $r_{\ell}\in\mathbb{R}^{n_{\ell}}$ in $\mathcal{O}(n_{\ell})$. Under these assumptions the computational cost of one Multigrid Monte Carlo update in Alg. 1 grows in proportion to the problem size $n_{L}$, as we will see in the following.
We start by analysing the cost of the random smoother in Alg. 3. Generating the random vector $\xi_{\ell}^{\mathrm{diag}}\in\mathbb{R}^{n_{\ell}}$ incurs a cost of $\mathcal{O}(n_{\ell})$ and generating $\xi_{\ell}^{\mathrm{LR}}\in\mathbb{R}^{\beta}$ a cost of $\mathcal{O}(\beta^{p_{\Gamma}})$ where $p_{\Gamma}=1$ if $\Gamma$ is diagonal and $p_{\Gamma}=2$ otherwise. Computation of $B_{\ell}\xi_{\ell}^{\text{LR}}$ in line 10 is $\mathcal{O}(\beta n_{\ell})$, while all other steps in the computation of $r_{\ell}$ and the tridiagonal solve for $\theta_{\ell}^{*}$ in line 11 are $\mathcal{O}(n_{\ell})$. Since the matrices $B_{\ell}$ and $B_{\ell}^{*}$ are both of size $n_{\ell}\times\beta$, the cost of the low-rank update in (3.14) is $\mathcal{O}(\beta n_{\ell})$.
@@
$\mathrm{Cost}_{\mathrm{other}}(\beta,n_{\ell})\leq C_{\mathrm{other}}(1+\beta^{p_{\Gamma}})n_{\ell}$ (3.17)
@@
[…]
@@
$\mathrm{Cost}_{\mathrm{MGMC}}(L)\leq C_{\mathrm{MG}}n_{L}\cdot\begin{cases}1+\beta&\text{if $\Gamma$ is diagonal,}\\ 1+\beta^{2}&\text{otherwise.}\end{cases}$ (3.20)
@@
$\mathrm{Cost}_{\mathrm{MGMC}}(\ell)\leq\begin{cases}C_{\mathrm{other}}(1+\beta^{p_{\Gamma}})n_{\ell}+\nu\mathrm{Cost}_{\mathrm{Gibbs}}(n_{\ell},\beta)+\gamma\mathrm{Cost}_{\mathrm{MGMC}}(\ell-1)&\text{for $\ell>0$}\\ \mathrm{Cost}_{\mathrm{coarse}}(\beta,\nu_{0},n_{0})&\text{for $\ell=0$}\end{cases}$
@@
$\mathrm{Cost}_{\mathrm{MGMC}}(L)\leq\sum_{\ell=1}^{L}\gamma^{L-\ell}\Big{(}C_{\mathrm{Gibbs}}^{(1)}\nu\beta^{p_{\Gamma}}+C_{\mathrm{MG}}^{(1)}(\beta,\nu)\cdot n_{\ell}\Big{)}+C_{\mathrm{MG}}^{(2)}(\beta,\nu_{0},n_{0})\gamma^{L}n_{0}$ (3.21)
@@
$C_{\mathrm{MG}}^{(1)}(\beta,\nu)$ $:=C_{\mathrm{other}}(1+\beta^{p_{\Gamma}})+C_{\mathrm{Gibbs}}^{(2)}\nu(1+\beta)$ (3.22)
@@
$C^{\prime}_{\mathrm{MG}}(\beta,\nu_{0},\nu,n_{0}):=C_{\mathrm{Gibbs}}^{(1)}\nu\beta^{p_{\Gamma}}n_{0}^{-1}+C_{\mathrm{MG}}^{(1)}(\beta,\nu)+C_{\mathrm{MG}}^{(2)}(\beta,\nu_{0},n_{0}).$ (3.24)
@@
The only two practically relevant recursion schemes are the V-cycle ($\gamma=1$) and the W-cycle ($\gamma=2$). Thus, when $A_{L}$ arises from the discretisation of a differential operator, the assumption in (3.19) is naturally satisfied. Consider for example a lowest order finite element discretisation on a uniform rectangular grid for which $\rho_{G}=2^{-d}\leq\frac{1}{4}$ in $d\geq 2$ dimensions. Then, $\gamma\rho_{G}\leq\frac{1}{2}<1$ for $\gamma\leq 2$.
In the above analysis we have neglected the costs of any setup operations, such as the computation of the Cholesky factorisation of $\widetilde{A}_{0}$ or the factorisation of $\Gamma$ which are required only once at the very beginning of a simulation. This is a valid assumption since these costs are amortised when a large number of samples is generated.
@@ 4 Theory
This section presents our main theoretical results. We first put the relationship between random samplers and linear solvers already mentioned in Section 3.3 into a wider context: in Section 4.1 we argue that according to the pioneering work in [Fox.C_Parker_2017_AcceleratedGibbsSampling], the convergence of the Multigrid Monte Carlo sampler is closely related to the convergence of deterministic multigrid solvers for linear systems of equations. This is made more explicit in Section 4.2, where we show that the Multigrid Monte Carlo update Alg. 1 leaves the target distribution invariant, and Section 4.3, where we prove grid-independent convergence of the algorithm. Throughout the discussion we highlight the connection with the theory for deterministic multigrid methods; to streamline the presentation we defer proofs to the appendices. Finally, in Section 4.4, we apply and extend the theory to conditional distributions and Bayesian inverse problems introduced in Section 2.2. We conclude with a result on the optimality of the MGMG sampler for drawing samples from the target distribution in Section 4.5.
4.1 Random sample splitting
Let $A=A^{\top}$ be a symmetric positive definite matrix. As pointed out in [Fox.C_Parker_2017_AcceleratedGibbsSampling], the problem of sampling from the multivariate normal distribution $\mathcal{N}(\mu,A^{-1})$ with mean $\mu=A^{-1}f$ and covariance $A^{-1}$ is closely related to the problem of finding the solution of a linear system with the precision matrix $A$. More specifically, assume that we want to generate samples $\theta\sim p(\theta)d\theta$ with the density

### v3

@@ 3.2 Random smoothers
A central ingredient of Alg. 1 is the $\mathrm{RandomSmoother}()$ procedure in lines 7 and 16, which can be seen as a generalisation of the Gibbs smoother. All random smoothers considered here are based on splitting a general symmetric matrix $A$ into two parts as A=:MNA=:M-N, where the matrix $M$ defines the splitting. The generic update procedure is written down in Alg. 2 where here and in the following we assume that the input $\theta$ is drawn from a multivariate normal distribution. As the following lemma shows, it is always possible to symmetrise a given random smoother defined by the splitting A=MNA=M-N.
Algorithm 2 Random Smoother for updating $\theta\mapsto\theta^{\prime}$ based on the matrix splitting A=MNA=M-N
1:procedure $\mathrm{RandomSmoother}$(AA, $M$, $f$, $\theta$)
@@ 3.3 Relationship to deterministic smoothers
would result in a deterministic smoother defined by the splitting A=MNA=M-N (see e.g. [undefaaq, section 4.2.2]). Under certain conditions on the splitting matrix $M$, repeated applications of the update in (3.5) can be used to iteratively solve the linear system $Au=f$. For example, when splitting the symmetric matrix $A$ into its diagonal $D$, strict lower triangular part $L$ and strict upper diagonal part $L^{\top}$
@@
$M$ $=M^{\mathrm{SOR}}=\frac{1}{\omega}D+L,$ $N$ $=N^{\mathrm{SOR}}=\frac{\omega-1}{\omega}D-L^{\top},$ (3.7)
where $\omega>0$ is referred to as relaxation parameter in the context of iterative solvers. As discussed extensively in [undef, undefae], there is a close relationship between classical smoothers for the solution of the linear system $Au=f$ and what we call random smoothers for sampling from the multivariate normal distribution $\mathcal{N}(\mu,A^{-1})$ with $\mu=A^{-1}f$ in Section 3.2. For example, the random pendant of the deterministic SOR smoother in (3.7) with $\omega=1$ is the Gibbs-sampler (see [undefae, Table 2]), which we will refer to as the Gibbs-smoother from now on to make the connection more explicit.
@@ 3.4 Random smoother with low-rank update for Bayesian inference
$\theta^{\prime}_{\ell}$ $=\theta_{\ell}^{*}-G_{\ell}(B_{\ell}^{\top}\theta_{\ell}^{*})$
@@
$G_{\ell}=(\omega^{-1}D_{\ell}+L_{\ell})^{-1}B_{\ell}\left(\Gamma+B_{\ell}^{\top}\left(\omega^{-1}D_{\ell}+L_{\ell}\right)^{-1}B_{\ell}\right)^{-1},$ (3.14)
which can be precomputed once at the beginning of the simulation. An equivalent backward Gibbs sampler with low rank correction can be obtained by swapping $L_{\ell}\leftrightarrow L_{\ell}^{\top}$ in (3.13) and (3.14). Putting everything together we arrive at Alg. 3. Obviously, $C_{\ell}$ and $G_{\ell}$ can be precomputed once and used in all subsequent calls, so that the setup costs are amortised for large numbers of samples.
@@
8:  Compute $G_{\ell}=C_{\ell}\left(\Gamma+B_{\ell}^{\top}C_{\ell}\right)^{-1}$
@@
$\theta^{\prime}_{\ell}=\theta_{\ell}^{*}-G_{\ell}(B_{\ell}^{\top}\theta_{\ell}^{*})$ (3.15)
@@ 3.5 Coarse level sampler
On the coarsest level we have to make a choice for the $\mathrm{CoarseSampler()}$ procedure that generates a new sample $\theta^{\prime}_{0}$ in line 3 of Alg. 1. Using Alg. 5, which consists of repeated applications of the random smoother in Alg. 2, corresponds to the approach in [undef]. It turns out that this choice is sufficient to guarantee that Alg. 1 leaves the target distribution invariant (see Section 4.2).
@@ 3.6 Cost analysis
In this section, we derive an upper bound on the cost of one MGMC update in Alg. 1 as a function of the problem size $n_{L}$ and the rank $\beta\ll n_{L}$ of the measurement operator $B_{L}$. We assume that this rank is fixed and independent of $L$. The value of $\beta$ depends on the application; in our numerical experiments in Section 5 we set $\beta=8$ in $d=2$ dimensions and $\beta=32$ in $d=3$. We analyse this cost under the assumption that the matrices $A_{\ell}$, $I_{\ell-1}^{\ell}$ and $I_{\ell}^{\ell-1}$ are sparse, with some upper bound on the number of entries per row, independent of the problem size. For the local basis functions in finite element spaces considered here this is naturally the case. Consequently, any matrix-vector products with $A_{\ell}$ or $I_{\ell-1}^{\ell}$ incur a cost of $\mathcal{O}(n_{\ell})$, in the same way as any tridiagonal solve $(\omega^{-1}D_{\ell}+L_{\ell})\theta_{\ell}=r_{\ell}$ or $(\omega^{-1}D_{\ell}+L_{\ell}^{\top})\theta_{\ell}=r_{\ell}$ for a single right-hand side $r_{\ell}\in\mathbb{R}^{n_{\ell}}$ in $\mathcal{O}(n_{\ell})$. Under these assumptions the computational cost of one Multigrid Monte Carlo update in Alg. 1 grows in proportion to the problem size $n_{L}$, as we will see in the following.
We start by analysing the cost of the random smoother in Alg. 3. Generating the random vector $\xi_{\ell}^{\mathrm{diag}}\in\mathbb{R}^{n_{\ell}}$ incurs a cost of $\mathcal{O}(n_{\ell})$ and generating $\xi_{\ell}^{\mathrm{LR}}\in\mathbb{R}^{\beta}$ a cost of $\mathcal{O}(\beta^{p_{\Gamma}})$ where $p_{\Gamma}=1$ if $\Gamma$ is diagonal and $p_{\Gamma}=2$ otherwise. Computation of $B_{\ell}\xi_{\ell}^{\text{LR}}$ in line 10 is $\mathcal{O}(\beta n_{\ell})$, while all other steps in the computation of $r_{\ell}$ and the tridiagonal solve for $\theta_{\ell}^{*}$ in line 11 are $\mathcal{O}(n_{\ell})$. Since the matrices $B_{\ell}$ and $G_{\ell}$ are both of size $n_{\ell}\times\beta$, the cost of the low-rank update in (3.15) is $\mathcal{O}(\beta n_{\ell})$.
@@
$\mathrm{Cost}_{\mathrm{other}}(\beta,n_{\ell})\leq C_{\mathrm{other}}(1+\beta)n_{\ell}$ (3.18)
@@
$n_{\ell-1}\,\leq\,\rho_{G}\,n_{\ell}\quad\text{and}\quad\rho_{G}\,\gamma\,<\,1\qquad\text{for\;\;$\gamma:=\max_{\ell=0,\dots,L}\{\gamma_{\ell}\}$}.$ (3.20)
@@
$\mathrm{Cost}_{\mathrm{MGMC}}(L)\leq C_{\mathrm{MG}}n_{L}\cdot\begin{cases}1+\beta&\text{if $\Gamma$ is diagonal}\\ \beta^{p_{\Gamma}}&\text{otherwise}.\end{cases}$ (3.21)
@@
$\mathrm{Cost}_{\mathrm{MGMC}}(\ell)\leq\begin{cases}C_{\mathrm{other}}(1+\beta)n_{\ell}+\nu\mathrm{Cost}_{\mathrm{Gibbs}}(n_{\ell},\beta)+\gamma\mathrm{Cost}_{\mathrm{MGMC}}(\ell-1)&\text{for $\ell>0$}\\ \mathrm{Cost}_{\mathrm{coarse}}(\beta,\nu_{0},n_{0})&\text{for $\ell=0$}\end{cases}$
@@
$\mathrm{Cost}_{\mathrm{MGMC}}(L)\leq\sum_{\ell=1}^{L}\gamma^{L-\ell}\Big(C_{\mathrm{Gibbs}}^{(1)}\nu\beta^{p_{\Gamma}}+(1+\beta)C_{\mathrm{MG}}^{(1)}(\nu)\cdot n_{\ell}\Big)+C_{\mathrm{MG}}^{(2)}(\beta,\nu_{0},n_{0})\gamma^{L}n_{0}$ (3.22)
@@
$C_{\mathrm{MG}}^{(1)}(\nu)$ $=C_{\mathrm{other}}+C_{\mathrm{Gibbs}}^{(2)}\nu$ (3.23)
@@
$C^{\prime}_{\mathrm{MG}}(\beta,\nu_{0},\nu,n_{0}):=C_{\mathrm{Gibbs}}^{(1)}\nu\beta^{p_{\Gamma}}n_{0}^{-1}+(1+\beta)C_{\mathrm{MG}}^{(1)}(\nu)+C_{\mathrm{MG}}^{(2)}(\beta,\nu_{0},n_{0}).$ (3.25)
@@
The only two practically relevant recursion schemes are the V-cycle ($\gamma=1$) and the W-cycle ($\gamma=2$). Thus, when $A_{L}$ arises from the discretisation of a differential operator, the assumption in (3.20) is naturally satisfied. Consider for example a lowest order FE discretisation on a uniform rectangular grid for which $\rho_{G}=2^{-d}\leq\frac{1}{4}$ in $d\geq 2$ dimensions. Then, $\gamma\rho_{G}\leq\frac{1}{2}<1$ for $\gamma\leq 2$.
Remark 3.4.
As will be discussed in Section 4, the parameters $\nu_{0}$, $\nu_{1}$, $\nu_{2}$ and $\gamma$ are independent of the number of levels $L$. For fixed values of these parameters, the convergence rate of the MGMC sampler is independent of $L$. Hence, the constant $C_{\mathrm{MG}}$ in (3.21) is independent of $L$ as well.
3.6.1 Setup costs
In addition to the above online costs that are incurred in every MGMC iteration, there are also offline costs for precomputing required quantities such as the matrices $C_{\ell}$ and $G_{\ell}$ in Alg. 3, as well as the Cholesky factorisation for the coarsest level. These setup costs are in general amortised when generating a large number of samples, but we quantify them here for completeness. In particular, the following theorem shows that the setup costs of Alg. 1 grow linearly in the problem size $n_{L}$.
Theorem 3.5 (MGMC setup cost).
Assume that (as in Theorem 3.2) the problem size $n_{\ell}$ decreases geometrically on the coarser levels, i.e. there is a constant $\rho_{G}<1$ such that
$n_{\ell-1}\,\leq\,\rho_{G}\,n_{\ell}\quad\text{for all $\ell=1,2,\dots,L$}.$ (3.26)
Then there exist constants $C_{\mathrm{MG}}^{(\mathrm{setup},1)}$ and $C_{\mathrm{MG}}^{(\mathrm{setup},2)}$, which only depend on $\rho_{G}$, such that the setup cost of Alg. 1 can be bounded by
$\mathrm{Cost}^{(\mathrm{setup})}_{\mathrm{MGMC}}(L)$ $\leq C_{\mathrm{MG}}^{(\mathrm{setup},1)}\beta^{2}n_{L}+C_{\mathrm{MG}}^{(\mathrm{setup},2)}\beta^{3}L+\mathrm{Cost}_{\mathrm{coarse}}^{(\mathrm{setup})}(\beta,n_{0}).$ (3.27)
Depending on the choice in line 3 of Alg. 1, the setup cost on the coarsest level is bounded by
[…]
for some constants $C_{\mathrm{Gibbs}}^{(\mathrm{setup},1)}$, $C_{\mathrm{Gibbs}}^{(\mathrm{setup},2)}$ and where $\mathrm{Cost}^{(\mathrm{setup})}_{\mathrm{Cholesky}}(n_{0})$ is the cost for computing the Cholesky factorisation of the $n_{0}\times n_{0}$ matrix $\widetilde{A}_{0}$.
Proof. See Appendix C.1.
Naively the setup cost of the coarse level Cholesky factorisation is $\mathcal{O}(n_{0}^{3})$, but the complexity will be reduced if the sparsity of the $n_{0}\times n_{0}$ matrix $\widetilde{A}_{0}$ can be exploited. However, since $n_{0}\ll n_{L}$, the setup cost on the coarsest level $\mathrm{Cost}_{\mathrm{coarse}}^{(\mathrm{setup})}(\beta,n_{0})$ is in general neligible.
3.6.2 Memory requirements
As the following theorem shows, the memory requirements of Alg. 1 also grow linearly in $n_{L}$.
Theorem 3.6 (MCMC memory requirements).
Assume that (as in Theorem 3.2) the problem sizes $n_{\ell}$ satisfy (3.26). Then there exists a constant $C_{\mathrm{MG}}^{(\mathrm{mem})}$, which only depends on $\rho_{G}$, such that the total memory requirement of Alg. 1 can be bounded by
$\mathrm{Memory}_{\mathrm{MGMC}}(L)$ $\leq C_{\mathrm{MG}}^{(\mathrm{mem})}(1+\beta)n_{L}+\beta L+\mathrm{Memory}_{\mathrm{coarse}}(\beta,n_{0}).$ (3.29)
Depending on the choice in line 3 of Alg. 1, the memory requirements on the coarsest level are bounded by
[…]
where $\mathrm{Memory}_{\mathrm{Chol}}(n_{0})$ is the memory required to store the Cholesky factorisation of $\widetilde{A}_{0}$.
Proof. See Appendix C.2.
The memory footprint of the Cholesky factorisation is $\mathcal{O}(n_{0}^{2})$. It can be reduced by exploiting the sparsity of $\mathcal{A}_{0}$. Because of $n_{0}\ll n_{L}$ the storage requirements on the coarsest level are negligible.
@@ 4 Theory
This section presents our main theoretical results. We first put the relationship between random samplers and linear solvers already mentioned in Section 3.3 into a wider context: in Section 4.1 we argue that according to the pioneering work in [undef, undefae], the convergence of the Multigrid Monte Carlo sampler is closely related to the convergence of deterministic multigrid solvers for linear systems of equations. This is made more explicit in Section 4.2, where we show that the Multigrid Monte Carlo update Alg. 1 leaves the target distribution invariant, and Section 4.3, where we prove grid-independent convergence of the algorithm. Throughout the discussion we highlight the connection with the theory for deterministic multigrid methods; to streamline the presentation we defer proofs to the appendices. Finally, in Section 4.4, we apply and extend the theory to conditional distributions and Bayesian inverse problems introduced in Section 2.2. We conclude with a result on the optimality of the MGMG sampler for drawing samples from the target distribution in Section 4.5.
4.1 Random sampling based on matrix splittings
Let $A=A^{\top}$ be a symmetric positive definite matrix. As pointed out in [undefae], the problem of sampling from the multivariate normal distribution $\mathcal{N}(\mu,A^{-1})$ with mean $\mu=A^{-1}f$ and covariance $A^{-1}$ is closely related to the problem of finding the solution of a linear system with the precision matrix $A$. More specifically, assume that we want to generate samples $\theta\sim p(\theta)d\theta$ with the density